OpenStudy (anonymous):

Find all solutions of the equation sin^2x=2sinx+3 The answer is A+Bkpi where k is any integer and 0

OpenStudy (anonymous):

I know that A is 3/2pi but I am having trouble finding B , I've tried putting 2 but apparently I'm wrong can someone help ?

OpenStudy (ranga):

Let y = sin(x). Put it in the given equation and you will have a quadratic equation. Solve for y and now replace y with sin(x)

OpenStudy (anonymous):

Got I'm going to try, thanks :)

OpenStudy (anonymous):

*it

OpenStudy (ranga):

yw

OpenStudy (anonymous):

So I've gotten y=3 and y=1

OpenStudy (anonymous):

so if I replace I should find where sinx=3 and sinx=1 ?

OpenStudy (ranga):

yes. If any of the quadratic solution is invalid discard it.

OpenStudy (anonymous):

Okay so y=1 doesn't work and so how do I find sinx=3 ?

OpenStudy (ranga):

sin(x) cannot be greater than 1. So discard sinx = 3. sin(x) = 1 is good.

OpenStudy (anonymous):

Okay so if sin(x)=1 then six(x) is pi/2 which the value 1/2 for B when I imput it into my online hw says its wrong, hmmmm I don't know what to do :/ thank you for your effort at least :)

OpenStudy (ranga):

No. If sin(x) = 1 then x = pi/2 is one solution. But remember the sine function repeats itself every 2pi. And so there will be countless x for which sin(x) will be 1: pi/2, pi/2 + 2pi, pi/2 + 4pi, ... So the solution for x = pi/2 + k(2pi) where k is an integer. Compare this to A + Bkpi A + Bkpi = pi/2 + 2kpi So A = pi/2 and B = 2

OpenStudy (anonymous):

Got it thank you so much !

OpenStudy (ranga):

you are welcome.